please help

Evaluate the integral. 2 (2x + 2)dx 3 2 Sizx+2)dx= (2x + 2)dx = (Simplify your answer.) 3
Answer #1

Given (2012) du (2n+2) dr = (2*2 +20) 2)+7+262) -(*)*+269)) = 474 -(9+6) =8-15 - -